OpenStudy (anonymous):

the perimeter of a standard size rug is 32 ft. the length is 2 ft longer than the width. what's the width? what's the length?

OpenStudy (anonymous):

2w+2l=32 l=w+2 2w+2(w+2)=32 there's a start

OpenStudy (campbell_st):

draw a picture  so 32 = 2(w + 2 + w) solve for w

OpenStudy (anonymous):

2w+2w+4=32 4w+4=32 -4 -4 4w = 28 4w =4w w=7

OpenStudy (anonymous):

w=7+2= 9 so width is 9 ?

OpenStudy (anonymous):

length is 9 width is 7

OpenStudy (anonymous):

oooh i get it thanks

OpenStudy (anonymous):

sure thing